# Target Specific De Novo Design of Drug Candidate Molecules with Graph Transformer-based Generative Adversarial Networks

### v6

@@ 2.10 Investigation of Binding Properties of the de novo Showcase Molecule
Figure 4: The output of the structural analysis of selected molecules: Capivasertib (the ligand from the crystal structure PDB id: “4GV1”) [42] and the showcase de novo generated molecule: “Pyrrolo[1,2-a]pyrimidin-4(1H)-one” derivative ($MOL\_01\_027820$ in Figure 3). Block I: Molecular docking results: (A) AKT1 crystal complex structure with the co-crystallized ligand: Capivasertib (binding free energy in docking: -9.517 kcal/mol), and (B) The best pose in the molecular docking of the showcase de novo generated molecule $MOL\_01\_027820$, to the structurally resolved binding site of AKT1 (binding free energy in docking: -9.686 kcal/mol). Gray dashed lines represent Van der Waals interactions. Blue lines represent hydrogen bonds. Yellow dashed lines represent salt bridges. Block II: The molecular dynamics simulation output. The occupancy values of: (C) the single binding pattern of capivasertib, (D) the first binding pattern of $MOL\_01\_027820$, and (E) the second binding pattern of $MOL\_01\_027820$ in complex with AKT1 obtained from the MD simulations. Noninteracting backbone atoms are hidden for visual clarity. Only the interactions above 10% occupancy values are reported. Domain names are noted on the figure as given in the KLIFS database [58]. Abbreviations: I-VII represents $\beta$-sheet numbers, g.l represents glycine-rich loop, c.l represents catalytic loop, GK represents gatekeeper residue, and xDFG represents highly conserved kinase residues, linker represents the loop that connects the hinge domain to $\alpha$C-helix. Gray dashed lines represent Van der Waals interactions. Blue lines represent hydrogen bonds and water bridges. Yellow dashed lines represent salt bridges. The cyan line represents orthogonal multipolar interaction. MD-derived occupancy values are shown on each interaction. Block III: RMSD plots of the compounds in MD simulations: (F) capivasertib and (G) $MOL\_01\_027820$ in each MD simulation. The illustration of rotation by RMSF plots of compounds: (H) capivasertib and (I) $MOL\_01\_027820$ for different binding patterns observed in MD simulations. For visual clarity, similar RMSF values on ligand atoms are calculated together.) as “active”.
In summary, we found that the binding features of capivasertib and the conformations of AKT-1 were conserved during the simulations. The simulation analysis results of $MOL\_01\_027820$ revealed two different binding patterns. Still, they kept the molecule within the binding cleft, conserved the interactions within the adenine region and conserved the protein’s general DFG-in and $\alpha$C chain-in conformation. This indicates that DrugGEN was able to learn the binding features of the target protein (Figure 4).
@@ 2.11 Attention map visualisation
Artificial neural networks are generally characterised as black box models that lack interpretability. On the other hand, transformer-based models weigh intrinsic pairwise relationships between data tokens through the attention mechanism, which can be utilised to visualise the important relationships in the data and, therefore, interpret what the model learns. Adopting this strategy, we visualised the average attention scores (produced by the transformer encoder module of the generator network) for each atom in the selected de novo generated molecules to assess critical atoms and bonds according to the model’s perspective. We observed the correspondence between the atoms that received high attention scores and those interacting with AKT1 protein’s binding pocket residues according to the molecular docking analysis (Figure 5 and Figure S7). A high correspondence would indicate that the model learned which atoms are critical for enabling physical interaction with the target.
In Figure 5, interactions between the atoms of de novo molecules and residues in the AKT1 binding pocket are visualised via 2-D protein-ligand interaction diagrams for three different de novo molecules (Figure 5A, C, and E), using a 4 Angstrom cutoff. Also, in Figure 5, the same de novo molecules’ atoms with the highest average attention scores are visualised with green, yellow and red, indicating the importance appointed by the model (Figure 5B, D, and F). The de novo molecules employed in this analysis were randomly selected from the 30 promising molecules given in Figure 3.
Considering all three cases in Figure 5, the model identifies most of the atoms interacting with the target residues by assigning high attention scores. For example, in Figure 5A and B, N+H atoms forming both salt bridges and hydrogen bonds with the residue Glu234, and HN and N atoms forming hydrogen bonds with the residues Asp274 and Ala230, respectively, are assigned the highest attention values (i.e., the green coloured ligand atoms). The same pattern is observed for other interactions in Figure 5. Overall, 20 out of 22 ( 91%) ligand atoms in pairwise interactions were detected by the generator network of DrugGEN. Looking at the same examples from the other angle, 20 out of 24 ( %83) ligand atoms with top attention scores were in contact with the target residues, thus playing a critical role in the interaction. Even though the model did not use any data regarding pairwise molecular interactions during training and validation, it still managed to learn substructures crucial for binding to AKT1 and designed molecules accordingly. These findings illustrate the utility of attention maps to understand the model’s generative process. We provided two additional examples in Figure S7 with a similar outcome.
@@ 3 Discussion and Conclusion
In this study, we developed the DrugGEN system to generate target-centric drug-like molecules automatically. DrugGEN combines the GAN architecture with graph transformers to create an end-to-end system that can design novel ligands given the target protein. DrugGEN is trained and evaluated to generate new ligands of the AKT1 protein. DrugGEN displayed a better performance compared to SOTA models in fundamental performance metrics, pointing to its high generation efficiency and capacity. Considering physicochemical metrics and structural/topological features, we showed that the DrugGEN model can generate de novo molecules with similar molecular characteristics to real inhibitors of the AKT1 protein. Further in silico analyses (i.e., molecular docking and MD simulations) were conducted to assess the interaction-related characteristics of de novo molecules, the results of which indicated their high potential in AKT1 targeting. Intending to present a tool to the community, we openly shared the code base, datasets, all results and trained models in our code repository at https://github.com/HUBioDataLab/DrugGEN and as a web-based tool with a graphical interface at https://huggingface.co/spaces/HUBioDataLab/DrugGEN where users can easily generate de novo molecules via employing the desired DrugGEN model. Our downstream analysis pipeline (Figure 1B and C) can be used to select the best drug candidate molecules from the large-scale output of any generative molecule design model in a hierarchical manner to be directed to the synthesis and wet-lab validation steps.
@@
One essential point is that, unlike many previous generative models, DrugGEN operates at large molecular sizes similar to actual drugs and drug-like bioactive compounds (Figure S8) at the maximum size of 45 heavy atoms so that its output can be useful under different contexts even though this choice elevated the complexity of the problem. At this point, an important question is what steps should be taken after obtaining the de novo molecule records from a generative model so that they could ultimately be utilised in real-world applications of drug discovery, biotechnology or material science. In this work, we touched on this topic and proposed an in silico downstream analysis pipeline (Figure 1B and C) for an early selection of promising molecules. Our pipeline can be employed to select the best drug candidate molecules from the large-scale output of any generative model in a hierarchical manner to be directed to the synthesis and wet-lab validation steps.
@@
Considering limitations, the adversarial training principle of GANs presented challenges, including significant divergence between the generator and discriminator losses, mode collapse, and the vanishing gradient [59, 60, 61]. However, adopting the Wasserstein GAN with gradient penalty (WGAN-GP) [62] has proven effective in mitigating these issues. This study employed random real molecular graphs as input to the generator network instead of noise. The utilisation of large molecular graphs (i.e., composed of 45 heavy atoms) elevates the complexity of the generative task, making it highly difficult for the model to find a mapping between the input noise and the high dimensional latent space of drug-like small molecules. On the other hand, learning the mapping between random small molecules and the ones that are binders of the selected target is considered a more approachable problem, which yielded high scores in the end.
@@
In further studies, we plan to design and train: (i) DrugGEN models for other target proteins in the human proteome; (ii) GAN models that utilise sequence-based transformers using a valid and robust molecular notation such as SELFIES [63] together with a successful molecular representation learning model, e.g., SELFormer [64]; (iii) models that utilise fragments/ sub-structures of molecules as its building blocks, instead of individual atoms and bonds, to be able to generate highly synthesisable de novo compounds. Additionally, we plan to incorporate target protein features into the molecule generation process directly -also adopted in conventional structure-based drug design- to yield more successful learning regarding molecular structural constraints. Finally, we plan to improve the molecular generation process by incorporating the high-level functional properties of real drugs and drug candidate molecules (along with their structural features) into the model training procedure in the context of heterogeneous biomedical knowledge graphs [65].This architecture is intended to facilitate understanding the relationships between the structural and functional properties of small molecules and thereby enhance the design process.
@@ 4.1 Datasets
Experimental bioactivities, i.e., quantitative measurements of physical interactions between drug-like compounds and their target proteins, are the second data type used in the training of DrugGEN. The bioactivity data was also retrieved from the ChEMBL database. We applied various filters for standardisation, such as target type: “single protein”, assay type: “binding assay”, standard type: “=”, and pChEMBL value: “not null” (i.e., roughly correspond to curated activity data points). Then, bioactivity data belonging to the AKT1 target protein were selected from the filtered bioactivity dataset. The finalised dataset contains ligand interactions of the human AKT1 (“CHEMBL4282”) protein with a pChEMBL value equal to or greater than 6 (i.e., IC50 <= 1 µM) as well as SMILES notations of these ligands. This activity dataset was extended by adding the drug molecules sourced from the DrugBank database [67] that are known to interact with human AKT1 protein. The filtering process excluded molecules larger than 45 heavy atoms and resulted in a dataset of 2,405 active/interacting small molecule ligands to train generative models tailored to AKT. Their heavy atom size distribution is given in Figure S8.
@@ 4.2 Graph Encodings (Featurization)
In DrugGEN, input molecules are represented as graphs, each composed of two parts: an annotation matrix (which contains information about the atom features) and an adjacency matrix (which includes information on the presence of atomic bonds and their features). The annotation and adjacency matrices of the compounds were generated using the RDKit [68] 67 library and the SMILES representation of molecules. The annotation matrix has a size of 45 by 9, representing the maximum length (the number of heavy atoms) of the molecule and nine types of atoms (i.e., C, O, N, F, S, P, Cl, I, and the null / no atoms case), respectively. These one-hot-encoded atom types are used as node features of the molecule graphs. The adjacency matrix is 45 by 45 by 5, displaying the bond information of the respective atoms of the molecule (cells of the one hot encoded third dimension of the tensor; 0th: no bond between the corresponding atoms, 1st: single, 2nd: double, 3rd: triple, and 4th: aromatic bond).
@@ 4.3 The Architecture of DrugGEN
The generator ($G$) module (Figure 1D) employs transformer encoder blocks [40] which operate on graph-based data [69]. For this, both the annotation and adjacency matrices are required to be processed in the same module (details regarding the dimensions and the context of annotation and adjacency matrices are given above, in the section entitled “Graph Encodings - Featurization”). $G$ takes randomly selected real drug-like molecules from the training dataset as input instead of random noise which is usually employed as the input of the generator modules of GANs. This approach helps the model effectively handle the high complexity and sparsity of molecular graphs (i.e., 45*9 + 45*45*5 = 18,630 elements to represent a molecule). The input is fed through individual multi-layered perceptrons (MLPs) for annotation and adjacency matrices, both of which consists of four layers (i.e., input: 16, hidden1: 64, hidden2: 64, and output: 128 dimensions). These MLPs are utilized to create embedings of annotation and adjacency matrices with $d_{k}$ (default: 128) dimensions. Afterward, the input is fed to the transformer encoder module. Here, firstly, input is sent to the self attention mechanism. In the classic transformer architecture, $Q$, $K$ and $V$ variables are the representations of the same input sequence. Attention is calculated by the scaled dot product of $Q$ and $K$, after that, the attention is multiplied by $V$ to form the final product [40]. In the graph transformer setting, $Q_{m}$, $K_{m}$ and $V_{m}$ are the variables representing the annotation matrix of the molecule. However, here, attention weights are calculated as the multiplication of the adjacency matrix ($A_{m}$) of the molecules with the scaled dot product of $Q_{m}$ and $K_{m}$. Then, attention weights are multiplied with $V_{m}$ to create the final representation of the annotation matrix. The new representation of the adjacency matrix is the concatenated version of the attention weights as described in the study by Dviwedi et al. (2020) and Vignac et al. (2022) [69, 70]. For our model, output dimension size of the transformer is 128 for both the annotation and adjacency. The calculation of the attention is formulated below:
$Attention(Q_{m},K_{m},V_{m})=softmax(\frac{Q_{m}K_{m}^{T}}{\sqrt{d_{k}}}A_{m})V_{m}$ (1)

### v7

@@ 2.8 Experimental validation of de novo molecules for AKT1 inhibition
Figure 4: Structural analysis of capivasertib (bound ligand in 4GV1) and five de novo generated molecules (Mol. ID 1–5) that were selected for experimental validation. (A) 2D visualisation of capivasertib and Molecules 1–5, with their unique identifiers assigned during molecule generation, QED and SA values. (B) 3D protein-ligand interactions of capivasertib and Molecules 1–5 after molecular docking with AKT1 (PDB ID: 4GV1). The interactions are visualised with active site residues of AKT1 highlighted in beige color. Different interaction types are depicted as follows: blue lines represent hydrogen bonds, yellow dashed lines indicate salt bridges, gray dashed lines denote hydrophobic interactions, orange dashed lines highlight $\pi$-cation interactions, and green lines represent halogen bonds. (C) Initial binding orientations of capivasertib, Molecules 1 and 2 at the starting point of molecular dynamics (MD) simulations. (D) Key protein-ligand interactions observed during MD simulations, visualised with interacting residues and interaction types. The depicted poses represent the most populated conformations from each simulation. (E) Root-mean-square deviation (RMSD) values of capivasertib, Molecules 1 and 2 in complex with AKT1. (F) Root-mean-square fluctuation (RMSF) values of ligand atoms in the same complexes. Abbreviations: Exp. Mol. ID: Experimental molecule identifier. I-VII represents $\beta$-sheet numbers, g.l represents glycine-rich loop, c.l represents catalytic loop, GK represents gatekeeper residue, and xDFG represents highly conserved kinase residues; linker represents the loop that connects the hinge domain to $\alpha$C-helix. Gray dashed lines represent Van der Waals interactions. Blue lines represent hydrogen bonds and water bridges. Green lines indicate halogen bonds. Yellow dashed lines represent salt bridges. Directional interactions were noted only if the occupancy value was found above 10%; however, for visual clarity, occupancy values of the water bridges were stated only if they were above 30%.
2.9 Investigation of the binding properties of molecules using molecular dynamics
Here, we performed molecular dynamics (MD) simulations and evaluated the binding characteristics of five synthesised de novo molecules with AKT1. The simulations were performed on the crystal’s co-crystalized ligand, capivasertib, as well as the docked poses of Molecules 1-5. The starting orientations of all these compounds were visualised within the ATP cofactor binding site in Figures 4C and S8.
MD simulations showed that all compounds remained stable (RMSD < 1.6 Å) (Figures 4E and S8), with no exceptional movements (i.e., low RMSF) in most cases (Figures 4F and S8). Capivasertib maintained near-constant hydrogen bonds with key residues (Tyr229 and Ala230) and additional water-mediated interactions with Lys158, Asn279, and Asp292 (Figure 4D). Molecule 1, which exhibited the most potent inhibition (IC50 = 1.89 µM), formed strong hydrogen bonds with hinge residue Ala230 (96%) and interacted with surrounding acidic residues, suggesting its activity derives from a similar interaction network as capivasertib—though its lack of additional hinge contacts and robust hydrogen bond networks may slightly limit its efficacy (Figure 4D). In contrast, Molecule 2 formed halogen bonds with hinge residues Tyr229 and Ala230 and hydrogen bonds with other critical residues (Lys179, Glu234, Glu278, Thr291, Asp292), accounting for its second most potent inhibitory effect (Figure 4D). Details about the MD results of synthesised molecules are provided in Supplementary Material (S10).
@@ 2.10 Attention map visualisation
We visualised the attention scores of DrugGEN’s generator transformer-encoder as an explainability analysis to evaluate critical atomic interactions in de novo molecules. Specifically, we assessed the correspondence between ligand atoms with high attention scores and those identified as critical in molecular docking (i.e., atoms directly interacting with AKT1 residues). As shown in Figure 5, DrugGEN assigned high attention scores to key atoms involved in hydrogen bonds and salt bridges—accurately identifying  91% of the interacting ligand atoms—demonstrating that, despite not being trained on explicit pairwise interaction data, the generator network effectively learned substructures essential for AKT1 binding, underscoring DrugGEN’s potential for target-centric molecule design. Additional information can be found in Supplementary Material (S13).
@@ 3 Discussion and Conclusion
In this study, we developed the DrugGEN system to generate target-centric drug-like molecules automatically. DrugGEN combines the GAN architecture with graph transformers to create an end-to-end system that can design novel ligands given the target protein. We further advanced the graph transformer architecture by incorporating a modified attention mechanism that more effectively integrates edge information, thereby reinforcing the representation of molecular structures. DrugGEN is trained and evaluated to generate new ligands of the AKT1 and CDK2 protein. DrugGEN displayed a better performance compared to SOTA models in fundamental performance metrics, pointing to its high generation efficiency and capacity. Considering physicochemical metrics and structural/topological features, we showed that the DrugGEN model can generate de novo molecules with similar molecular characteristics to real inhibitors of the AKT1 protein. Further in silico analyses (i.e., molecular docking and MD simulations) and in vitro enzymatic assays were conducted to assess the interaction-related characteristics and potency of the de novo molecules. The results of enzymatic assays indicated two out of five compounds have high potency at low micromolar concentrations, 1.89 µM (Molecule 1) and 48.6 µM (Molecule 2). Intending to present a tool to the community, we openly shared the code base, datasets, all results and trained models in our code repository at https://github.com/HUBioDataLab/DrugGEN and as a web-based tool with a graphical interface at https://huggingface.co/spaces/HUBioDataLab/DrugGEN where users can easily generate de novo molecules via employing the desired DrugGEN model.
@@
One essential point about DrugGEN is that, unlike many previous generative models, DrugGEN operates at large molecular sizes similar to actual drugs and drug-like bioactive compounds (Figure S11) at the mean and maximum number of 33 and 45 heavy atoms, respectively, so that its output can be useful under different drug development-related contexts even though this choice elevated the complexity of the modelling problem.
A current and essential question in the literature is what steps should be taken after obtaining the de novo molecule records from a generative model so that they could ultimately be utilised in real-world applications of drug discovery, biotechnology or material science. In this work, we explored this topic and introduced a downstream analysis pipeline—incorporating both in silico and in vitro analyses (Figure 1C)—to facilitate the early selection of promising molecules. Our hierarchical selection pipeline can be applied to the large-scale outputs of any generative model, enabling the identification of the most promising drug candidate molecules and forwarding them for subsequent (pre)clinical development steps.
There are studies in the literature that aim to solve the target-specific and drug-like molecule generation problem via multi-objective optimisation. Several of these studies perform this task based on predictions from external models [62, 63, 64, 65]. For example, FSM-DDTR leverages the predicted bioactivity (pIC50) values of molecules to construct a predictive model, which is then used as a reward function in the molecular generation module [64]. However, performing optimisation based on predicted values carries the risk of error propagation. Our model simplifies the process by removing intermediate prediction-centric steps and directly mapping random molecule distributions to those of the selected protein’s inhibitors, possibly reducing the chance of error propagation. Additionally, since our model does not rely on multiple explicit optimisation steps, the complexity of training and inference processes is potentially lower. This, in turn, facilitates the practical development and implementation of our models. In future studies, these aspects can be assessed.
@@
Considering limitations, first of all, the adversarial training principle of GANs presented challenges, including significant divergence between the generator and discriminator losses, mode collapse, and the vanishing gradient [66, 67, 68]. However, adopting the Wasserstein GAN with gradient penalty (WGAN-GP) [69] has partially mitigated these issues. Second, as previously stated, this study uses random real molecular graphs, rather than noise, as input to the generator network. Given that large molecular graphs (composed of up to 45 heavy atoms) significantly increase the complexity of the generative task, it is exceedingly difficult for the model to establish a robust mapping between pure noise and the high-dimensional latent space of drug-like molecules. By substituting real molecular graphs for noise, we help the model more effectively learn a mapping from these initial structures to the de novo inhibitor molecules of the selected protein. Notably, this design does not impose explicit constraints on molecular novelty. Instead, the WGAN-GP framework encourages the generation of chemically valid yet novel molecules by penalising unrealistic structures rather than enforcing similarity to the training set. As demonstrated in Table 1, the model effectively generates novel compounds while maintaining chemical validity and target specificity.
@@
In further studies, we plan to design and train: (i) DrugGEN models for other target proteins in the human proteome; (ii) GAN models that utilise sequence-based transformers using a valid and robust molecular notation such as SELFIES [70]together with a successful molecular representation learning model, e.g., SELFormer [71]; (iii) models that utilise fragments/ sub-structures of molecules as its building blocks, instead of individual atoms and bonds, to be able to generate highly synthesisable de novo compounds. Another interesting direction could be incorporating constraint optimisation techniques into DrugGEN’s architecture to better control the generation of molecules with desired properties. Such constraints could be integrated into the GAN’s loss function to guide the generation process toward molecules that meet specific physicochemical criteria, target binding requirements, or synthesizability thresholds. This enhancement would allow for more precise control over the generated molecules’ properties while maintaining DrugGEN’s current advantages in handling large molecular structures.
Additionally, we plan to incorporate target protein features into the molecule generation process directly—also adopted in conventional structure-based drug design—to yield more successful learning regarding molecular structural constraints. Finally, we plan to improve the molecular generation process by incorporating the high-level functional properties of real drugs and drug candidate molecules into the model training procedure in the context of heterogeneous biomedical knowledge graphs [72]. This architecture is intended to facilitate understanding the relationships between the structural and functional properties of small molecules and thereby enhance the design process.
@@ 4.1 Datasets
Experimental bioactivities, i.e., quantitative measurements of physical interactions between drug-like compounds and their target proteins, are the second data type used in the training of DrugGEN. The bioactivity data was also retrieved from the ChEMBL database. We applied various filters for standardisation, such as target type: “single protein”, assay type: “binding assay”, standard type: “=”, and pChEMBL value: “not null” (i.e., roughly correspond to curated activity data points). Then, bioactivity data belonging to the AKT1 target protein were selected from the filtered bioactivity dataset. The finalised dataset contains ligand interactions of the human AKT1 (“CHEMBL4282”) protein with a pChEMBL value equal to or greater than 6 (i.e., IC50 <= 1 µM) as well as SMILES notations of these ligands. This activity dataset was extended by adding the drug molecules sourced from the DrugBank database [74] that are known to interact with human AKT1 protein. The filtering process excluded molecules larger than 45 heavy atoms and resulted in a dataset of 2,405 active/interacting small molecule ligands to train generative models tailored to AKT. Their heavy atom size distribution is given in Figure S10. The same procedure was applied for the human CDK2 (CHEMBL301) protein, resulting in 1,817 compounds.
@@ 4.2 Graph Encodings (Featurization)
In DrugGEN, input molecules are represented as graphs, each composed of two parts: an annotation matrix (which contains information about the atom features) and an adjacency matrix (which includes information on the presence of atomic bonds and their features). The annotation and adjacency matrices of the compounds were generated using the RDKit [75] 67 library and the SMILES representation of molecules. The annotation matrix has a size of 45 by 9, representing the maximum length (the number of heavy atoms) of the molecule and nine types of atoms (i.e., C, O, N, F, S, P, Cl, I, and the null / no atoms case), respectively. These one-hot-encoded atom types are used as node features of the molecule graphs. The adjacency matrix is 45 by 45 by 5, displaying the bond information of the respective atoms of the molecule (cells of the one hot encoded third dimension of the tensor; 0th: no bond between the corresponding atoms, 1st: single, 2nd: double, 3rd: triple, and 4th: aromatic bond). For CDK2-targeting molecules, the annotation matrix has dimensions 38 by 10, with 38 denoting the maximum number of heavy atoms and 10 corresponding to the atom types (C, N, O, F, P, S, Cl, Br, I, and a null token). The adjacency matrix for these molecules is 38 by 38 by 5, with the third dimension similarly representing bond types as described for the AKT1 case.
@@ 4.3 The Architecture of DrugGEN
The generator ($G$) module (Figure 1D) employs transformer encoder blocks [41] which operate on graph-based data [76], a method that has been shown to be highly effective for molecular data analysis [43, 44, 45, 76]. For this, both the annotation and adjacency matrices are required to be processed in the same module (details regarding the dimensions and the context of annotation and adjacency matrices are given above, in the section entitled “Graph Encodings - Featurization”). $G$ takes randomly selected real drug-like molecules from the training dataset as input instead of random noise which is usually employed as the input of the generator modules of GANs. This approach aligns with studies in the literature that use real samples as input for GAN generators instead of random noise [46, 47] This approach helps the model effectively handle the high complexity and sparsity of molecular graphs (i.e., 45*9 + 45*45*5 = 18,630 elements to represent a molecule). The input is fed through individual multi-layered perceptrons (MLPs) for annotation and adjacency matrices, both of which consists of four layers (i.e., input: 16, hidden1: 64, hidden2: 64, and output: 128 dimensions). These MLPs are utilized to create embedings of annotation and adjacency matrices with $d_{k}$ (default: 128) dimensions. Afterwards, the input is fed to the transformer encoder module. In the classic transformer architecture [41], self-attention is computed by deriving query ($QKV$, applying a softmax, and finally multiplying by $V$ to form the final output. Graph transformers extend this mechanism to graph-structured data by incorporating the molecule’s adjacency matrix $A_{m}$ to account for the connectivity between nodes (atoms). In this setting, $Q_{m}$, Km and $V_{m}$ are derived from the annotation matrix of the molecule. Unlike the original transformer, the attention weights here are obtained by multiplying the scaled dot product of $Q_{m}$ and $K_{m}$ with the $A_{m}$. The resulting attention is then multiplied by $V_{m}$ to create the updated representation of the annotation matrix, while the updated representation of the adjacency matrix is formed through the concatenation of these attention weights as described in the study by Dviwedi et al. (2020) and Vignac et al. (2022) [76, 77]. The purpose of performing element-wise multiplication of the attention with the adjacency matrix is to enhance the attention scores by incorporating bond information into them. The attention matrix aims to identify how much each atom attends to other atoms in a pairwise manner, and this is based on both short-range (direct) and long-range (indirect) relationships/interactions between the atoms in the molecule. Therefore, attention values can be enhanced by introducing readily available bond information between the corresponding atoms (i.e., whether an atomic bond exists or not, and if so, what specific kind of bond). In our model, we begin with the graph transformer’s standard attention formulation. Our key modification is an additional multiplication step involving a shifted adjacency matrix ($A_{m}+1$). Specifically, we first multiply the attention by $A_{m}$ and then multiply by ($A_{m}+1$) to amplify the edge signal, thereby enhancing the model’s ability to capture subtle structural features. The calculation of the final attention is formulated below:
$Attention(Q_{m},K_{m},V_{m})=softmax(\frac{Q_{m}K_{m}^{T}}{\sqrt{d_{k}}}A_{m}(A_{m}+1))V_{m}$ (1)